# Free Descriptions of Convex Sets

### v2

@@ Appendix B Deriving the Presentation Degree from Extendability
    We begin carrying out the first stage of the above program and identify two simple types of relations. The first source for relations between generators in $\mathbb{V}_{d}$ arises from relations over $\mathsf{G}_{d}$. Indeed, if $\sum_{i}g_{i}x_{i}=0$ for $g_{i}\in\mathsf{G}_{d}$ and $x_{i}\in\mathbb{V}_{d}$ then $\sum_{i}gg_{i}x_{i}=0$ for any $g\in\mathsf{G}_{n}$. Such relations are always satisfied by the images $A_{n_{0}}x_{i}$. A second source for such relations arises from subgroups of $\mathsf{G}_{n}$ acting trivially on $\mathbb{V}_{d}$, which are presicely the centralizing subgroups of Definition 2.4. Centralizing subgroups yield a second source for relations, namely, the relations $(h-\mathrm{id})x=0$ for all $x\in\mathbb{V}_{d}$ and $h\in\mathsf{H}_{n,d}$. Thus, if $A_{n_{0}}$ extends to a morphism then $A_{n_{0}}(\mathbb{V}_{d})\subseteq\bigcap_{n\geq d}\mathbb{U}_{n}^{\mathsf{H}_{n,d}}$. Rather than attempt to enforce these constraints computationally, we make a standard simplifying assumption from the representation stability literature. Specifically, we assume that $\mathbb{U}_{d}$ is fixed by $\mathsf{H}_{n,d}$ (or a subgroup of it, see below) for all $n\geq d$, in which case there is a simple sufficient condition for the above constraints that we can impose computationally. This is precisely the assumption that $\mathscr{U}$ is a $\mathscr{V}$-module as in Definition 2.5. This terminology comes from a categorical approach to representation stability, see Appendix A. If $\mathscr{U}$ is a $\mathscr{V}$-module, then imposing $A_{n_{0}}(\mathbb{V}_{d})\subseteq\mathbb{U}_{d}$ is sufficient to guarantee $A_{n_{0}}(\mathbb{V}_{d})\subseteq\bigcap_{n\geq d}\mathbb{U}_{n}^{\mathsf{H}_{n,d}}$ since $\mathbb{U}_{d}$ is contained in the right-hand side of this inclusion. Imposing this sufficient condition can done computationally, as we do in Section 6.1. This concludes the first stage.
@@
    If =$\mathscr{F}=\{\mathbb{F}_{n}\}$ is a $\mathscr{V}$-module and $\{A_{n}\colon\mathbb{F}_{n}\to\mathbb{U}_{n}\}$ is a surjective morphism, then for any $n\geq d$ we have $\mathbb{U}_{n}=A_{n}(\mathbb{F}_{n})=A_{n}(\mathbb{R}[\mathsf{G}_{n}]\mathbb{F}_{d})=\mathbb{R}[\mathsf{G}_{n}]A_{n}(\mathbb{F}_{d})=\mathbb{R}[\mathsf{G}_{n}]A_{d}(\mathbb{F}_{d})=\mathbb{R}[\mathsf{G}_{n}]\mathbb{U}_{d},$ where we used the fact that $\mathscr{F}$ is generated in degree $d$; the equivariance of $A_{n}$; the fact that $A_{n}|_{\mathbb{F}_{d}}=A_{d}$ since $\{A_{n}\}$ is a morphism; and the surjectivity of $A_{d}$. This shows $\mathscr{U}$ is generated in degree $d$.
@@
    Appendix C Schur-Horn Orbitopes in AF Algebras
[…]
    The spectral theorem for $C^{*}$ algebras gives an (isometric) isomorphism of algebras between continuous functions on the spectrum $\sigma(x)=\{\lambda\in\mathbb{C}:x-\lambda I\textrm{ not invertible}\}$ of any self-adjoint $x\in\overline{\mathbb{A}_{\infty}}$ and the closed subalgebra it generates [90, Thm. 11.19]. This allows us to apply continuous functions $f\colon\sigma(x)\to\mathbb{C}$ to $x$ itself to obtain $f(x)\in\overline{\mathbb{A}_{\infty}}$. Using this functional calculus, for each self-adjoint $x\in\overline{\mathbb{A}_{\infty}}$ we get the positive unital linear functional on such functions sending $f\mapsto\tau(f(x))$. By the Riesz representation theorem [91, Thm. 6.19], there is a unique probability measure $\mu_{x}^{\tau}$ on $\sigma(x)$ satisfying $\mu_{x}^{\tau}(f)=\tau(f(x))$ for all continuous $f\colon\sigma(x)\to\mathbb{C}$, which is called the spectral measure of $x$ with respect to $\tau$. Finally, if $\sigma(x)=\{\lambda_{1},\ldots,\lambda_{q}\}$ is finite then we have a spectral decomposition $x=\sum_{i=1}^{q}\lambda_{i}p_{i}$ where $\{p_{i}\}$ are orthogonal projectors satisfying $p_{i}p_{j}=\delta_{i,j}p_{i}$, $\sum_{i}p_{i}=1$, and $\mu_{x}^{\tau}=\sum_{i=1}^{q}\tau(p_{i})\delta_{\lambda_{i}}$.
@@ Appendix C Schur-Horn Orbitopes in AF Algebras
    $\varphi_{n}^{*}(x)=\frac{1}{2}\sum_{i=1}^{2}(I\otimes e_{i})^{\top}x(I\otimes e_{i}),\qquad\textrm{ hence }\|\varphi_{n}^{*}x\|\leq\|x\|,$ (34)
@@
    To obtain conic descriptions for the above orbitopes, let $\mathcal{K}_{n}=\{xx^{*}:x\in\mathbb{A}_{n}\}\subseteq\mathbb{V}_{n}$ be the cone of positive-semidefinite Hermitian matrices in $\mathbb{A}_{n}$, so that $\mathcal{K}_{n}\subseteq K_{n+1}$ and $\overline{\mathcal{K}_{\infty}}=\{xx^{*}:x\in\overline{\mathbb{A}_{\infty}}\}$. We also have $\mathcal{P}_{n}(K_{n+1})=\mathcal{K}_{n}$ by (34), hence $\mathcal{P}_{n}(\overline{\mathcal{K}_{\infty}})=\mathcal{K}_{n}$. Therefore, the finite-dimensional description of Schur-Horn orbitopes [70, Eq. (19)] reads
@@
    Let $\lambda_{1},\ldots,\lambda_{q}$ are distinct real numbers, and let $m_{1},\ldots,m_{q}\in\mathbb{N}$ with sum $m=\sum_{i}m_{i}$. Then $\overline{\mathrm{SH}(x_{1})_{\infty}}$ with $x_{1}=\mathrm{diag}(\widetilde{\lambda})$ as above is
@@
    and $u_{n}=(0,0,-I,-m_{1}/m,\ldots,-m_{q}/m)$. Observing that $\{A_{n}\},\{A_{n}^{*}\},\{B_{n}\},\{B_{n}^{*}\}$ are all morphisms, that $u_{n}=u_{n+1}$ (under our embeddings), the descriptions in (35) are free and satisfy the hypohteses of Proposition 3.2(a). Furthermore, putting the norm $\|(p_{1},\ldots,p_{q})\|=\max_{i}\|p_{i}\|$ on $W_{\infty}$ and $\|(p_{1},\ldots,p_{q},y_{1},y_{2},v)\|\leq\max\{\|p_{i}\|,\|y_{j}\|,\|v\|_{\infty}\}$ on $\mathbb{U}_{\infty}$, we have $\overline{W_{\infty}}=\overline{\mathbb{V}_{\infty}}^{\oplus q}$ and $\overline{\mathbb{U}_{\infty}}=\overline{W_{\infty}}\oplus\overline{\mathbb{V}_{\infty}}^{\oplus 2}\oplus\mathbb{R}^{q}$. Moreover, $\|A_{n}\|_{\mathrm{op}}=1$, $\|B_{n}\|_{\mathrm{op}}\leq\max\{\sum_{i}|\lambda_{i}|,q\}$ for all $n$. Thus, both $A_{n}$ and $B_{n}$ extend to the continuous limit and Theorem 3.6 indeed applies and yields (36). ∎
    The Schur-Horn orbitopes $\mathrm{SH}(x_{1})_{n}$ are precisely the Schur-Horn orbitopes $\mathrm{SH}(\lambda)_{n}$ in (15) from Section 4.3 if $\mathbb{A}_{n}=\mathbb{R}^{m2^{n}\times m2^{n}}$ with $\mathsf{G}_{n}=\mathrm{O}(m2^{n})$. They also reduce to the permutahedra $\mathrm{SH}(x_{1})_{n}=\mathrm{Perm}(\lambda)_{n}$ as in (12) from Section 4.3 if $\mathbb{A}_{n}$ is the algebra of diagonal matrices in $\mathbb{M}_{n}$ with $\mathsf{G}_{n}=\mathrm{S}_{m2^{n}}$.
    Proposition C.2 yields a generalization of the Schur-Horn theorem to AF algebras. Indeed, let $\mathbb{D}_{n}=\{x\in\mathbb{V}_{n}:x\textrm{ diagonal}\}$ which is a unital subalgebra of $\mathbb{A}_{n}$ embedded in $\mathbb{D}_{n+1}$ under $\varphi_{n}$. For $x_{1}\in\mathbb{D}_{1}$, define
@@
    The sequence of linear maps $(\mathrm{diag}_{n}\colon\mathbb{V}_{n}\to\mathbb{D}_{n})_{n}$ extracting the diagonal of a Hermitian matrix in $\mathbb{V}_{n}$ extend to a bounded linear map $\mathrm{diag}\colon\overline{\mathbb{V}_{\infty}}\to\overline{\mathbb{D}_{\infty}}$ since $\mathrm{diag}_{n+1}\circ\varphi_{n}=\varphi_{n}\circ\mathrm{diag}_{n}$ and $\|\mathrm{diag}_{n}\|_{\mathrm{op}}=1$. The finite-dimensional Schur-Horn theorem states that $\mathrm{diag}_{n}(\mathrm{SH}(x_{1})_{n})=\mathrm{Perm}(x_{1})_{n}$. Using Proposition C.2, we obtain the following infinite-dimensional extension.
@@
    =$=\overline{\left\{\sum\nolimits_{i=1}^{q}\lambda_{i}p_{i}:p_{i}\in\overline{\mathcal{K}_{\infty}}\cap\overline{\mathbb{D}_{\infty}},\ \sum\nolimits_{i}p_{i}=1,\ \tau(p_{i})=m_{i}/m\right\}}$
@@
    The finite-dimensional Schur-Horn theorem yields $\mathrm{Perm}(x_{1})_{\infty}=\mathrm{diag}(\mathrm{SH}(x_{1})_{\infty})$, and taking closures yields a weaker statement since it involves the closure of the image of diag. Proposition C.3 shows that this closure can be removed.
    There is a considerable literature on extending the Schur-Horn theorem to infinite-dimensional setting, including the extension in [92] for operators with a finite spectrum, and in [93] for von Neumann algebras. As explained in [93, §1], removing the closure over the image of diag has been a major challenge in these more general settings. Our descriptions of limiting convex sets from Theorem 3.6 can be seen as resolving this challenge in our simpler setting of AF algebras.

### v3

@@ Appendix B Deriving the Presentation Degree from Extendability
    We begin carrying out the first stage of the above program and identify two simple types of relations. The first source for relations between generators in $\mathbb{V}_{d}$ arises from the action of $\mathsf{G}_{d}$ on $\mathbb{V}_{d}$ itself, without involving any higher dimensions. Indeed, if $\sum_{i}g_{i}x_{i}=0$ for $g_{i}\in\mathsf{G}_{d}$ and $x_{i}\in\mathbb{V}_{d}$ is such a relation then $\sum_{i}g_{i}A_{n_{0}}x_{i}=0$ so the image $A_{n_{0}}x_{i}$ satisfies the same relation. Moreover, the above relations in $\mathbb{V}_{d}$ give rise to relations in $\mathbb{V}_{n}$ for $n>d$ since if $\sum_{i}g_{i}x_{i}=0$ as above then $\sum_{i}gg_{i}x_{i}=0$ in $\mathbb{V}_{n}$ for any $g\in\mathsf{G}_{n}$. Once again, these higher-dimensional relations are always satisfied by $A_{n_{0}}x_{i}$. A second source for such relations arises from subgroups of $\mathsf{G}_{n}$ acting trivially on $\mathbb{V}_{d}$, which are precisely the centralizing subgroups of Definition 2.4. Centralizing subgroups yield a second source for relations, namely, the relations $(h-\mathrm{id})x=0$ for all $x\in\mathbb{V}_{d}$ and $h\in\mathsf{H}_{n,d}$. Thus, if $A_{n_{0}}$ extends to a morphism then $A_{n_{0}}(\mathbb{V}_{d})\subseteq\bigcap_{n\geq d}\mathbb{U}_{n}^{\mathsf{H}_{n,d}}$. Rather than attempt to enforce these constraints computationally, we make a standard simplifying assumption from the representation stability literature. Specifically, we assume that $\mathbb{U}_{d}$ is fixed by $\mathsf{H}_{n,d}$ (or a subgroup of it, see below) for all $n\geq d$, in which case there is a simple sufficient condition for the above constraints that we can impose computationally. This is precisely the assumption that $\mathscr{U}$ is a $\mathscr{V}$-module as in Definition 2.5. This terminology comes from a categorical approach to representation stability, see Appendix A. If $\mathscr{U}$ is a $\mathscr{V}$-module, then imposing $A_{n_{0}}(\mathbb{V}_{d})\subseteq\mathbb{U}_{d}$ is sufficient to guarantee $A_{n_{0}}(\mathbb{V}_{d})\subseteq\bigcap_{n\geq d}\mathbb{U}_{n}^{\mathsf{H}_{n,d}}$ since $\mathbb{U}_{d}$ is contained in the right-hand side of this inclusion. Imposing this sufficient condition can done computationally, as we do in Section 6.1. This concludes the first stage.
@@
    If $\mathscr{F}=\{\mathbb{F}_{n}\}$ is a $\mathscr{V}$-module and $\{A_{n}\colon\mathbb{F}_{n}\to\mathbb{U}_{n}\}$ is a surjective morphism, then for any $n\geq d$ we have $\mathbb{U}_{n}=A_{n}(\mathbb{F}_{n})=A_{n}(\mathbb{R}[\mathsf{G}_{n}]\mathbb{F}_{d})=\mathbb{R}[\mathsf{G}_{n}]A_{n}(\mathbb{F}_{d})=\mathbb{R}[\mathsf{G}_{n}]A_{d}(\mathbb{F}_{d})=\mathbb{R}[\mathsf{G}_{n}]\mathbb{U}_{d},$ where we used the fact that $\mathscr{F}$ is generated in degree $d$; the equivariance of $A_{n}$; the fact that $A_{n}|_{\mathbb{F}_{d}}=A_{d}$ since $\{A_{n}\}$ is a morphism; and the surjectivity of $A_{d}$. This shows $\mathscr{U}$ is generated in degree $d$.
@@
    Appendix C Schur–Horn Orbitopes in AF Algebras
    The permutahedra and Schur–Horn orbitopes studied in Section 4.3 are special cases of a more general construction arising in the field of operator algebras, and which is naturally analyzed using our framework. In this appendix, we present this unified view. Let $\mathbb{A}_{n}\subseteq\mathbb{M}_{n}\coloneqq\mathbb{C}^{m2^{n}\times m2^{n}}$ be a unital $\mathbb{C}$-subalgebra with the faithful tracial state $\tau_{n}(x)=\mathrm{Tr}(x)/m2^{n}$ which induces the inner product $\langle x,y\rangle_{n}=\tau_{n}(x^{*}y)$, and the max-singular value operator norm $\|\cdot\|_{(n)}$. Let $\varphi_{n}\colon\mathbb{M}_{n}\hookrightarrow\mathbb{M}_{n+1}$ be the unital algebra embedding $\varphi_{n}(x)=x\otimes I_{2}$, and note that $\|x\otimes I_{2}\|_{(n+1)}=\|x\|_{(n)}$ and $\tau_{n+1}\circ\varphi_{n}=\tau_{n}$, so $\varphi_{n}$ is an isometry with respect to the above inner product. Thus, both the operator norm $\|\cdot\|$ and the normalized trace $\tau$ do not depend on $n$ and extend to the limit $\mathbb{A}_{\infty}=\bigcup_{n}\mathbb{A}_{n}$, and we omit their subscripts above. Let $\overline{\mathbb{A}_{\infty}}$ be the closure of $\mathbb{A}_{\infty}$ with respect to the operator norm, which is now a $C^{*}$ algebra called an approximately finite-dimensional (AF) algebra [88, Chap. 3]. Note that $\tau$ extends continuously to $\overline{\mathbb{A}_{\infty}}$ since $|\tau(x)\|\leq\|x\|$ for all $x\in\mathbb{A}_{\infty}$, and hence $\tau$ defines a tracial state on $\overline{\mathbb{A}_{\infty}}$.
    The spectral theorem for $C^{*}$ algebras gives an (isometric) isomorphism of algebras between continuous functions on the spectrum $\sigma(x)=\{\lambda\in\mathbb{C}:x-\lambda I\textrm{ not invertible}\}$ of any self-adjoint $x\in\overline{\mathbb{A}_{\infty}}$ and the closed subalgebra it generates [89, Thm. 11.19]. This allows us to apply continuous functions $f\colon\sigma(x)\to\mathbb{C}$ to $x$ itself to obtain $f(x)\in\overline{\mathbb{A}_{\infty}}$. Using this functional calculus, for each self-adjoint $x\in\overline{\mathbb{A}_{\infty}}$ we get the positive unital linear functional on such functions sending $f\mapsto\tau(f(x))$. By the Riesz representation theorem [90, Thm. 6.19], there is a unique probability measure $\mu_{x}^{\tau}$ on $\sigma(x)$ satisfying $\mu_{x}^{\tau}(f)=\tau(f(x))$ for all continuous $f\colon\sigma(x)\to\mathbb{C}$, which is called the spectral measure of $x$ with respect to $\tau$. Finally, if $\sigma(x)=\{\lambda_{1},\ldots,\lambda_{q}\}$ is finite then we have a spectral decomposition $x=\sum_{i=1}^{q}\lambda_{i}p_{i}$ where $\{p_{i}\}$ are orthogonal projectors satisfying $p_{i}p_{j}=\delta_{i,j}p_{i}$, $\sum_{i}p_{i}=1$, and $\mu_{x}^{\tau}=\sum_{i=1}^{q}\tau(p_{i})\delta_{\lambda_{i}}$.
@@ Appendix C Schur–Horn Orbitopes in AF Algebras
[…]
@@
    To obtain conic descriptions for the above orbitopes, let $\mathcal{K}_{n}=\{xx^{*}:x\in\mathbb{A}_{n}\}\subseteq\mathbb{V}_{n}$ be the cone of positive-semidefinite Hermitian matrices in $\mathbb{A}_{n}$, so that $\mathcal{K}_{n}\subseteq K_{n+1}$ and $\overline{\mathcal{K}_{\infty}}=\{xx^{*}:x\in\overline{\mathbb{A}_{\infty}}\}$. We also have $\mathcal{P}_{n}(K_{n+1})=\mathcal{K}_{n}$ by (34), hence $\mathcal{P}_{n}(\overline{\mathcal{K}_{\infty}})=\mathcal{K}_{n}$. Therefore, the finite-dimensional description of Schur–Horn orbitopes [68, Eq. (19)] reads
@@
[…]
@@
    and $u_{n}=(0,0,-I,-m_{1}/m,\ldots,-m_{q}/m)$. Observing that $\{A_{n}\},\{A_{n}^{*}\},\{B_{n}\},\{B_{n}^{*}\}$ are all morphisms, that $u_{n}=u_{n+1}$ (under our embeddings), the descriptions in (35) are dimension-free and satisfy the hypotheses of Proposition 3.2(a). Furthermore, putting the norm $\|(p_{1},\ldots,p_{q})\|=\max_{i}\|p_{i}\|$ on $W_{\infty}$ and $\|(p_{1},\ldots,p_{q},y_{1},y_{2},v)\|\leq\max\{\|p_{i}\|,\|y_{j}\|,\|v\|_{\infty}\}$ on $\mathbb{U}_{\infty}$, we have $\overline{W_{\infty}}=\overline{\mathbb{V}_{\infty}}^{\oplus q}$ and $\overline{\mathbb{U}_{\infty}}=\overline{W_{\infty}}\oplus\overline{\mathbb{V}_{\infty}}^{\oplus 2}\oplus\mathbb{R}^{q}$. Moreover, $\|A_{n}\|_{\mathrm{op}}=1$, $\|B_{n}\|_{\mathrm{op}}\leq\max\{\sum_{i}|\lambda_{i}|,q\}$ for all $n$. Thus, both $A_{n}$ and $B_{n}$ extend to the continuous limit and Theorem 3.6 indeed applies and yields (36). ∎
    The Schur–Horn orbitopes $\mathrm{SH}(x_{1})_{n}$ are precisely the Schur–Horn orbitopes $\mathrm{SH}(\lambda)_{n}$ in (15) from Section 4.3 if $\mathbb{A}_{n}=\mathbb{R}^{m2^{n}\times m2^{n}}$ with $\mathsf{G}_{n}=\mathrm{O}(m2^{n})$. They also reduce to the permutahedra $\mathrm{SH}(x_{1})_{n}=\mathrm{Perm}(\lambda)_{n}$ as in (12) from Section 4.3 if $\mathbb{A}_{n}$ is the algebra of diagonal matrices in $\mathbb{M}_{n}$ with $\mathsf{G}_{n}=\mathrm{S}_{m2^{n}}$.
    Proposition C.2 yields a generalization of the Schur–Horn theorem to AF algebras. Indeed, let $\mathbb{D}_{n}=\{x\in\mathbb{V}_{n}:x\textrm{ diagonal}\}$ which is a unital subalgebra of $\mathbb{A}_{n}$ embedded in $\mathbb{D}_{n+1}$ under $\varphi_{n}$. For $x_{1}\in\mathbb{D}_{1}$, define
@@
    The sequence of linear maps $(\mathrm{diag}_{n}\colon\mathbb{V}_{n}\to\mathbb{D}_{n})_{n}$ extracting the diagonal of a Hermitian matrix in $\mathbb{V}_{n}$ extend to a bounded linear map $\mathrm{diag}\colon\overline{\mathbb{V}_{\infty}}\to\overline{\mathbb{D}_{\infty}}$ since $\mathrm{diag}_{n+1}\circ\varphi_{n}=\varphi_{n}\circ\mathrm{diag}_{n}$ and $\|\mathrm{diag}_{n}\|_{\mathrm{op}}=1$. The finite-dimensional Schur–Horn theorem states that $\mathrm{diag}_{n}(\mathrm{SH}(x_{1})_{n})=\mathrm{Perm}(x_{1})_{n}$. Using Proposition C.2, we obtain the following infinite-dimensional extension.
@@
    $=\overline{\left\{\sum\nolimits_{i=1}^{q}\lambda_{i}p_{i}:p_{i}\in\overline{\mathcal{K}_{\infty}}\cap\overline{\mathbb{D}_{\infty}},\ \sum\nolimits_{i}p_{i}=1,\ \tau(p_{i})=m_{i}/m\right\}}$
@@
    The finite-dimensional Schur–Horn theorem yields $\mathrm{Perm}(x_{1})_{\infty}=\mathrm{diag}(\mathrm{SH}(x_{1})_{\infty})$, and taking closures yields a weaker statement since it involves the closure of the image of diag. Proposition C.3 shows that this closure can be removed.
    There is a considerable literature on extending the Schur–Horn theorem to infinite-dimensional setting, including the extension in [91] for operators with a finite spectrum, and in [92] for von Neumann algebras. As explained in [92, §1], removing the closure over the image of diag has been a major challenge in these more general settings. Our descriptions of limiting convex sets from Theorem 3.6 can be seen as resolving this challenge in our simpler setting of AF algebras.